If a surface can provide a friction of 5N and we apply a force of 10N then it starts moving but how friction can stip its motion when applied force is more than the friction force
1
Expert's answer
2018-06-06T10:44:08-0400
When applied force is more than the friction force friction can stop car’s motion:
F_net=ma=F-F_f=10-5=5 N>0 N.
So,
a>0.